OpenStudy (anonymous):

Write an Equation in slope intercept, point slope, or standard form for the line with the given information. Explain why you chose the form you used. A. passes through (-1,4) and (-5,2) B. Slope 2 intercept -4 C. has an x intercept of 6 and a y intercept of 3 D. Passes through (1,2) with slope -5/3

OpenStudy (anonymous):

Do you know how to find the slope between 2 points?

OpenStudy (anonymous):

Can you explain it please?

OpenStudy (anonymous):

Subtract the y values and x values, then divide the remaining y/x

OpenStudy (anonymous):

So for A. (-1,4) and (-5,2) (2-4) and -5-(-1) so -2/-4=....1/2 so the slope for the points (-1,4) and (-5,2) is 1/2

OpenStudy (anonymous):

Now pick any of the points, either (-1,4) or (-5,2) and using the slope that you just found, write an equation for the line. You will use the point slope formula.

OpenStudy (anonymous):

So I'll pick the point (-1,4) and with the slope 1/2 write the following formula y-4=(1/2)(x-(-1).....y=4=.5(x+1)....This will be the answer for A

OpenStudy (anonymous):

Thank You So Much!